# Early stages of collective cell invasion: Biomechanics

### v1

@@ 2.1.1. Durotaxis
The interplay between strain and substrate stiffness creates a feedback loop that governs cell dynamics and collective behavior. When cells adhere to a deformable substrate, they exert traction forces that deform it. The deformation changes the substrate’s stiffness, which in turn influences cell migration and adhesion dynamics. This feedback loop is governed by the strain-stiffening relationship in Eqs (7)-(8). The strain-induced changes in stiffness affect the Hamiltonian strain term through Eq (9). A more elaborated CPM model that explains durotaxis from cell-ECM forces and focal adhesion dynamics is presented in Ref. ren20 .
@@
Fig 1: Simulated active and passive cells. Dynamics of red cells (passive, $E_{\theta c}=1$ kPa) and blue cells (active, embedded into a disk of passive cells, $E_{\theta m}=15$ kPa), with substrate strains (black arrows). Initial condition: circle-shaped distribution with 10 randomly placed blue cells. Panels show different iterations: (a) t=50, (b) t=500, (c) t=1000. See Video S1. Line segments indicate strain magnitude and orientation above a given threshold vanOers .
@@
By means of Eqs (1)-(9), we have simulated the evolution of a circular agglomerate of red passive non-active cells with embedded blue active cells placed in random positions. The only differences among cells are their stiffness thresholds: $E_{\theta c}=1$ kPa (passive cells) and $E_{\theta m}=15$ kPa (active cells). Non-active cells maintain equilibrium under substrate rigidity, while active cells respond to strain gradients generated by passive cells. The result is that active cells move towards the borders of the agglomerate and may leave it, as shown in Fig 1. For further details, see Video S1.
2.1.2. Incorporation of Migration Forces in the Model
To achieve directed migration of the active cells, we introduce a global long range active migration force $\mathbf{f}_{m}$. Long range force transmission of mechanical origin is responsible for phenomena such as waves in antagonistic migration assays rod17 , stress fibres mechanics oak17 , collective cell durotaxis sun16 , etc. The migration force must be incorporated into the finite element part of the model by modifying the Navier equation:
$\mathbf{K}\mathbf{u}=\mathbf{f}+\mathbf{f}_{m}.$ (10a)
Here, we propose an idealized simple long range migration force: $\mathbf{f}_{m}$ is zero for passive cells, whereas it depends on the distance to a point of reference (push force) or to a point of attraction (pull force) for active cells. A push migration force on node $i$ of the active cells is
$\mathbf{f}_{mi}=\mu\mathbf{d}_{mi}=\mu(\mathbf{d}_{m}-\mathbf{d}_{i}),$ (10b)
where $\mathbf{d}_{m}$ is the position of a reference point opposite to the target location, and $\mathbf{d}_{i}$ is the position of node $i$ of the active cell. The force $\mathbf{f}_{m}$ points towards $d_{m}$, away from the target point, and modifies the displacement $\mathbf{u}$, whose discrete gradient for an extension towards the target point is $\mathbf{0}-\mathbf{f}_{mi}$. In turn, this generates strains $\mathbf{\varepsilon}$ in Eq (7) whose magnitude increases with $|\mathbf{d}_{m}-\mathbf{d}_{i}|$. These strains enter the CPM through the durotaxis Eq. (6) and push the cell towards the target location (under cellular extension towards it). As an alternative to Eq. (10b), the pull migration force is
$\mathbf{f}_{mi}=\mu\mathbf{d}_{ia}=\mu(\mathbf{d}_{i}-\mathbf{d}_{a}),$ (10c)
where $\mathbf{d}_{a}$ is the position of the target location (attraction point). The force $\mathbf{f}_{mi}$ of Eq. (10c) also points away from the target location $\mathbf{d}_{a}$ and pulls the cell towards the attraction point. The difference between push and pull forces is that the push force is stronger the further the distance to $\mathbf{d}_{m}$ is (and therefore the shorter the distance to $\mathbf{d}_{a}$ is), whereas the pull force is weaker the shorter the distance to $\mathbf{d}_{a}$ is. The push force provides a more efficient attraction to the target point and therefore it is our force of choice.
Fig 2: Simulated migration force. Migration of active blue cells ($E_{\theta m}=15$ kPa) and passive red cells ($E_{\theta c}=1$ kPa). Substrate strains are shown by black arrows. Blue cells are attracted to the lower left corner. The initial condition is the same as in Fig 1. Panels at iterations: (a) t=1000, (b) t=1500, (c) t=3000. See Video S2.
The active force modifies the behavior of the active cells, directing them preferentially towards the point of attraction and away from the reference point: The strain gradient created by the migration force directs the movement of the cells. Figure 2 shows how the active blue cells respond to the migration force by moving towards the attraction center whereas the passive red cells are insensitive to it. The black arrows represent the strains in the substrate, which are influenced by the migration force. These simulations illustrate the influence of substrate stiffness, active forces, and strain feedback on collective cell behavior. These mechanisms are crucial in tissue morphogenesis and the self-organization of cells into functional structures. Next, we specify different adhesion properties and threshold stiffnesses for cells of different phenotypes that are responsible for their diverse behavior.
2.2. Influence of Cell-Cell and Cell-Substrate Adhesion
@@
To incorporate a hybrid cell type that exhibits both active behavior and a tendency to adhere to similar cells, we modify the cell-cell and cell-substrate adhesion parameters, $J(\sigma(\mathbf{x}),\sigma(\mathbf{x}^{\prime}))$ in Eq (5). In this subsection, we analyze how these parameters influence the interactions between various cell types and the substrate, leveraging adhesion values from Graner ; Glazier to identify distinct cellular phases. While phenotypes may result from cell biochemical processes (which we discuss in a forthcoming paper), we need to assign different mechanical properties to different cell types. In a similar vein, different mechanical properties are assigned to leader and follower (tip and stalk) cells resulting from Notch signaling dynamics in angiogenesis veg20 .
2.2.1. Pattern Types According to Adhesion Parameters
@@
We consider two cell types (passive and active). Their adhesion dynamics depend on the stiffness thresholds ($E_{\theta c}$ for passive cells, $E_{\theta m}$ for active cells), the cell populations ($N_{m}$ for active cells, $N_{c}$ for passive cells, and $N_{\text{total}}$ overall) and the cell-cell and cell-substrate adhesion parameters, $J(\sigma(\mathbf{x}),\sigma(\mathbf{x}^{\prime}))$ in Eq (5). The latter adopt different values with respect to a reference adhesion $J_{\text{ref}}$, which we denote by
@@
  • $J_{cm}=j_{cm}J_{\text{ref}}$: passive cell-active cell adhesion,
  • $J_{mm}=j_{mm}J_{\text{ref}}$: active cell-active cell adhesion,
  • $J_{ms}=j_{ms}J_{\text{ref}}$: active cell-substrate adhesion.
Lower (resp. higher) adhesion parameter values facilitate (resp. impede) adhesion. The relative values between adhesion parameters are critical: If $j_{cs}=j_{cm}>j_{cc}$ for example, passive cells prefer to adhere to each other preferably than adhering to the substrate or to active cells. Through a qualitative analysis, we have identified in the Supplementary Table S1 seven dynamic patterns based on variations in adhesion parameters, substrate stiffness, and cell populations:
@@
    Clusters of passive and active cells sharing boundaries, each cell preferring their own type: When $j_{cc}<j_{mm}\approx j_{cm}\approx j_{cs}\approx j_{ms}$, cells cluster with their own type, and secondarily with the opposite type, over sticking to the substrate. See Fig. 3(a).
@@
    Separate clusters consisting of only passive or active cells with no inter-species adhesion: Strong intra-species adhesion ($j_{qq}$, $q=c,m$) and weak inter-species adhesion, $j_{cm}\gg j_{qq}$, promote segregation. See Fig. 3(b).
@@
    Separated passive and active cells, with unbound active cells: Active cells exhibit weak interactions, while passive cells cluster tightly. Flexible substrates enhance active cell freedom. For structured separations, this pattern requires $j_{cm},j_{mm}\gg j_{ms}$ and balanced populations, $N_{m}\approx N_{c}$. See Fig. 3(c).
@@
    Disordered compact pattern with active cells forming the outer rim: Cells of different species cluster together when inter-species adhesion dominates, $j_{cm}\ll j_{mm}$. See Fig. 3(d).
@@
    Active cells push passive cells apart yielding a fragmented pattern: active cells exert pressure, disrupting passive cell cohesion. Differences in stiffness thresholds, $E_{\theta c}\neq E_{\theta m}$, and low $j_{ms},j_{cm}$ drive this pattern. See Fig. 3(e).
@@
    Single active cells escape from a cluster of passive cells, active cells are at the periphery and moving away: Active cells surround passive cell groups when $j_{ms}\ll j_{mm},j_{cm}$ and $N_{m}\ll N_{c}$. Passive cells act as the central reference group. See Fig. 3(f).
@@
    Cluster metastasis: active cells form clusters around passive cells: Active cells group together in a ‘metastasis’ configuration around passive cells when $j_{mm}\leq j_{ms}\ll j_{cm}$. See Fig. 3(g).
@@
Fig 3: Changes in adhesion parameters. A total of $N_{total}=111$ cells are randomly placed inside a circular enclosure, with $N_{m}$ active cells (in blue) and passive adhesive cells (red). (a) $E_{\theta c}=E_{\theta m}=1$kPa, $j_{cs}=1$, $j_{cc}=1/2$, $j_{cm}=2$, $j_{ms}=1$, $j_{mm}=2$, $N_{m}=56$. (b) same as (a) except $j_{cm}=11$. (c) $E_{\theta c}=1$ kPa, $E_{\theta m}=15$ kPa, $j_{cs}=1$, $j_{cc}=1$, $j_{cm}=4$, $j_{ms}=1/4$, $j_{mm}=6$, $N_{m}=56$. (d) same as (a) except $j_{mm}=10$. (e) $E_{\theta c}=1$ kPa, $E_{\theta m}=15$ kPa, $j_{cs}=1$, $j_{cc}=1/2$, $j_{cm}=2$, $j_{ms}=1$, $j_{mm}=12$, $N_{m}=10$. (f) $E_{\theta c}=1$ kPa, $E_{\theta m}=15$ kPa, $j_{cs}=1$, $j_{cc}=1$, $j_{cm}=4$, $j_{ms}=1/4$, $j_{mm}=6$, $N_{m}=10$. (g) same as (f) except $j_{cm}=10$, $j_{ms}=1/2$, $j_{mm}=1/2$. (h) same as (a) except $j_{mm}=3$. (i) same as (e) except $j_{cm}=10$, $j_{ms}=1/4$. At $t=3000$, the corresponding pattern types are: (a) 1, (b) 2, (c) 3, (d) 4, (e) 5, (f) 6, (g) 7, (h) 1,2,3, and (i) 3,6.
@@
Figure 3 illustrates these patterns: It shows cellular configurations of types 1 to 7 as well as mixtures of different types in Figs. 3(h) and 3(i). To ensure reproducibility of our results, Supplementary Table S1 has 21 configurations and lists the conditions under which different patterns are found. For example, entry # 24 in the table combines configurations 2, 5, 7, 9, 21. It corresponds to an aggregate of 111 cells, of which 10 (a much smaller number) are active (configurations 2,5). The thresholds for stiffness (the Young modulus) in Eq. (9) are $E_{\theta c}=1$ kPa (passive cells) and $E_{\theta m}=15$ kPa (active cells) as in configuration 7. Thus, active cells move more easily than passive ones. The relative adhesion parameters are $j_{ms}=1/4$, $j_{cs}=j_{cc}=j_{mm}=1$, $j_{cm}=10$, which correspond to configurations 9, 21. The adhesion parameter between active cells and substrate is smaller than that between passive cells and substrate and active-active and passive-passive cells, whereas the adhesion parameter between active and passive cells is much higher. With these values, active cells tend to disassociate themselves from passive cells whereas adhesions between like cell types or with the substrate are much more favorable. The resulting pattern is number 6, corresponding to Fig. 3(f). Mixtures of patterns occur with intermediate parameter values, reflecting locally dependent choices by cells; see Supplementary Table S1.
2.2.2. Implementation of the Migration Force
Incorporating the active migration force to the CPM means solving Eqs. (2)-(10). Then the forces $\mathbf{f}$ and $\mathbf{f}_{m}$ contribute equally to the overall displacement vector $\mathbf{u}$ according to Eq. (10a). This approach fails to generate sufficient active displacement when the Hamiltonian variation from strains generated by substrate forces (stiffness force plus migration force) is less important than that generated by contact interactions with other cells and the substrate (which depends on adhesion parameter values). In other words, migration becomes negligible for some configurations of adhesion parameters when the Hamiltonian variation due to stiffness is greater than that of the migration force.

### v2

@@ III Young modulus threshold and adhesion in durotaxis
The interplay between strain and substrate stiffness creates a feedback loop that governs cell dynamics and collective behavior. When cells adhere to a deformable substrate, they exert traction forces that deform it. The deformation changes the substrate’s stiffness, which in turn influences cell migration and adhesion dynamics. This feedback loop is governed by the strain-stiffening relationship in Eqs (7)-(8). The strain-induced changes in stiffness affect the Hamiltonian strain term through Eqs. (6) and (9). In extension, we use the YMT of the target pixel $\mathbf{x}^{\prime}$, whereas we use the YMT of the source pixel $\mathbf{x}$ in retraction. Cells with a large YMT produce large strains in extension and move more frequently according to Eq. (8)-(9). Motile cells exhibit weak interactions, while passive cells with a smaller YMT cluster tightly. A more elaborated CPM explains durotaxis from cell-ECM forces and focal adhesion dynamics ren20 .
@@
Fig 1: Dynamics of motile and passive cells having different YMTs $E_{\theta}$. Initially, 10 blue motile cells ($E_{\theta m}=15$ kPa) are randomly embedded into a disk of 101 passive red cells ($E_{\theta c}=1$ kPa) with relative adhesion values $j_{cc}=j_{cm}=j_{mm}=2$, $j_{cs}=j_{ms}=1$, where $J_{kl}=j_{kl}J_{\text{ref}}$ (see Subsection III.2). Black segments indicate strain magnitude and orientation above a given threshold vanOers . Monte Carlo time steps (MCTS): (a) t=30 , (b) t=250, (c) t=500. See Video S1 suppl .
III.1 Effects of the Young modulus threshold
@@
By means of Eqs (1)-(9), we have simulated the evolution of a circular agglomerate of cells with different YMTs: $E_{\theta c}=1$ kPa (red non-motile cells) and $E_{\theta m}=15$ kPa (blue motile cells). Non-motile cells maintain equilibrium under substrate rigidity, while motile cells respond to strain gradients generated by cells of the other type. The result is that motile cells proceed towards the borders of the agglomerate, as shown in Fig 1. The aggregate breaks down in different ways depending on the initial distribution of cells. See Video S1 suppl .
III.2 Patterns due to adhesion and YMT
@@
To incorporate a hybrid cell type that exhibits both motile behavior and a tendency to adhere to similar cells, we modify the cell-cell and cell-substrate adhesion parameters, $J(\sigma(\mathbf{x}),\sigma(\mathbf{x}^{\prime}))$, in Eq (5). Their values influence the interactions between various cell types and the substrate, thereby allowing to identify cellular phases with distinct mechanical properties as phenotypes. The latter may result from cell biochemical processes (which we discuss in a forthcoming paper). In a similar vein, different mechanical properties are assigned to leader and follower (tip and stalk) cells resulting from Notch signaling dynamics in angiogenesis veg20 .
@@
We consider two cell types (passive and motile). Their adhesion dynamics depend on their YMT ($E_{\theta c}$ for passive, $E_{\theta m}$ for motile cells), the cell populations ($N_{m}$ motile, $N_{c}$ passive cells with $N_{\text{total}}=N_{c}+N_{m}$), and the cell-cell and cell-substrate adhesion parameters, $J(\sigma(\mathbf{x}),\sigma(\mathbf{x}^{\prime}))$, in Eq (5). The latter adopt different values with respect to a reference adhesion $J_{\text{ref}}$, which we denote by
@@
  • $J_{cm}=j_{cm}J_{\text{ref}}$: passive cell-motile cell adhesion,
  • $J_{mm}=j_{mm}J_{\text{ref}}$: motile cell-motile cell adhesion,
  • $J_{ms}=j_{ms}J_{\text{ref}}$: motile cell-substrate adhesion.
See Fig. 1. Lower (resp. higher) adhesion parameter values facilitate (resp. impede) adhesion. The relative values between adhesion parameters are critical: If $j_{cs}=j_{cm}>j_{cc}$ for example, passive cells prefer to adhere to each other preferably than adhering to the substrate or to motile cells. Through a qualitative analysis, we have identified in Appendix A seven dynamic patterns based on variations in adhesion parameters, substrate stiffness, and cell populations:
@@
    Clusters of passive and motile cells sharing boundaries, each cell preferring their own type: When $j_{cc}<j_{mm}\approx j_{cm}\approx j_{cs}\approx j_{ms}$, cells prefer clustering with their own type, and secondarily with the opposite type, over sticking to the substrate. See Fig. 2(a).
@@
    Separate clusters consisting of only passive or motile cells with no inter-species adhesion: Strong intra-species adhesion ($j_{qq}$, $q=c,m$) and weak inter-species adhesion, $j_{cm}\gg j_{qq}$, promote segregation. See Fig. 2(b).
@@
    Separated passive and motile cells, some single motile cells: Motile cells exhibit weak interactions, while passive cells cluster tightly. Flexible substrates enhance freedom of motile cells. For structured separations, this pattern requires $j_{cm},j_{mm}\gg j_{ms}$ and balanced populations, $N_{m}\approx N_{c}$. See Fig. 2(c).
@@
    Disordered compact pattern with motile cells forming the outer rim: Cells of different phenotype cluster together when inter-phenoype adhesion dominates, $j_{cm}\ll j_{mm}$. See Fig. 2(d).
@@
    Motile cells push passive cells apart yielding a fragmented pattern: Motile cells exert pressure, disrupting cohesion of passive cells. Differences in stiffness thresholds, $E_{\theta c}\neq E_{\theta m}$, and $j_{ms}<j_{cm}<j_{mm}$ drive this pattern. See Fig. 2(e).
@@
    Single motile cells escape from a cluster of passive cells, motile cells are at the periphery and moving away: Motile cells surround passive cell groups when $j_{ms}\ll j_{mm},j_{cm}$ and $N_{m}\ll N_{c}$. Passive cells act as the central reference group. See Fig. 2(f).
@@
    Cluster metastasis: Motile cells form clusters around passive cells: Motile cells group together in a ‘metastasis’ configuration around passive cells when $j_{mm}\leq j_{ms}\ll j_{cm}$. See Fig. 2(g).
@@
Fig 2: Changes in adhesion parameters produce different patterns. A total of $N_{total}=111$ cells are randomly placed inside a circular enclosure, with $N_{m}$ motile cells (blue) and passive adhesive cells (red) having $j_{cs}=1$ and $E_{\theta c}=E_{\theta m}=1$ kPa in Panels (a), (b), (d), (h), and $E_{\theta c}=1$ kPa, $E_{\theta m}=15$ kPa in Panels (c), (e), (f), (g), (i). Relative adhesion parameters are ($J_{kl}=j_{kl}J_{ref}$): (a) $j_{cc}=1/2$, $j_{cm}=2$, $j_{ms}=1$, $j_{mm}=2$, $N_{m}=56$. (b) same as (a) except $j_{cm}=11$. (c) $j_{cc}=1$, $j_{cm}=4$, $j_{ms}=1/4$, $j_{mm}=6$, $N_{m}=56$. (d) same as (a) except $j_{mm}=10$. (e) $j_{cc}=1/2$, $j_{cm}=2$, $j_{ms}=1$, $j_{mm}=12$, $N_{m}=10$. (f) $j_{cc}=j_{cs}=1$, $j_{cm}=4$, $j_{ms}=1/4$, $j_{mm}=6$, $N_{m}=10$. (g) same as (f) except $j_{cm}=10$, $j_{ms}=1/2$, $j_{mm}=1/2$. (h) same as (a) except $j_{mm}=3$. (i) same as (e) except $j_{cm}=10$, $j_{ms}=1/4$. At $t=3000$ MCTS, the corresponding pattern types are: (a) 1, (b) 2, (c) 3, (d) 4, (e) 5, (f) 6, (g) 7, (h) 1,2,3, and (i) 3,6.
@@
Figure 2 illustrates these patterns. Figs. 2(a), 2(b), 2(d), and 2(h) have equal YMTs in Eq. (9), $E_{\theta c}=E_{\theta m}=1$kPa, whereas Figs. 2(c), 2(e), 2(f), and 2(g) have different YMT, $E_{\theta c}=1$ kPa, $E_{\theta m}=15$ kPa. Only cells with different YMT show motile cells able to travel, thereby describing possible metastatic scenarios. Figure 2 shows cellular configurations of the main patterns 1 to 7 as well as mixtures of different patterns in Figs. 2(h) and 2(i). To ensure reproducibility of our results, Appendix A enumerates 21 configurations and lists in a table the conditions under which different patterns are found. To understand the table, take for example entry # 24 that combines configurations 2, 5, 7, 9, 21. It corresponds to an aggregate of 111 cells, of which 10 (a much smaller number) are motile (configurations 2,5) having different YMTs as in configuration 7. Thus, motile cells advance more easily than passive ones. The relative adhesion parameters are $j_{ms}=1/4$, $j_{cs}=j_{cc}=j_{mm}=1$, $j_{cm}=10$, which correspond to configurations 9, 21. The adhesion parameter between motile cells and substrate is smaller than that between passive cells and substrate and motile-motile and passive-passive cells, whereas the adhesion parameter between motile and passive cells is much higher. With these values, motile cells tend to disassociate themselves from passive cells whereas adhesions between like cell types or with the substrate are much more favorable. The resulting pattern is number 6 corresponding to Fig. 2(f). Mixtures of patterns occur with intermediate parameter values, reflecting locally dependent choices by cells; see Appendix A.
IV Active migration forces
To achieve directed migration of the motile cells, we introduce a global long range active migration force $\mathbf{f}_{m}$. Long range force transmission of mechanical origin is responsible for phenomena such as waves in antagonistic migration assays rod17 , stress fibres mechanics oak17 , collective cell durotaxis sun16 , etc. The migration force should be incorporated into the finite element part of the model by modifying the Navier equation:
$\mathbf{K}\mathbf{u}=\mathbf{f}+\mathbf{f}_{m}.$ (10a)
Here, we propose an idealized simple long range migration force: $\mathbf{f}_{m}$ is zero for passive cells, whereas it depends on the distance to a point of reference (push force) or to a point of attraction (pull force) for motile cells. A push migration force on node $i$ of the motile cells is
$\mathbf{f}_{mi}=\mu\mathbf{d}_{mi}=\mu(\mathbf{d}_{m}-\mathbf{d}_{i}),$ (10b)
where $\mathbf{d}_{m}$ is the position of a reference point opposite to the target location, and $\mathbf{d}_{i}$ is the position of node $i$ of the motile cell. We have kept the same tension per unit length as in Eq. (1), which is calibrated to the observed values of the total cell traction vanOers . Raising or lowering the value of $\mu$ in Eq. (10b) with respect to that in Table 1 does not change the outcome of the numerical simulations (see below). The force $\mathbf{f}_{m}$ points towards $d_{m}$, away from the target point, and modifies the displacement $\mathbf{u}$, whose discrete gradient for an extension towards the target point is $\mathbf{0}-\mathbf{f}_{mi}$. In turn, this generates strains $\mathbf{\varepsilon}$ in Eq (7) whose magnitude increases with $|\mathbf{d}_{m}-\mathbf{d}_{i}|$. These strains enter the CPM through the durotaxis Eq. (6) and push the cell towards the target location (under cellular extension towards it). As an alternative to Eq. (10b), the pull migration force is
$\mathbf{f}_{mi}=\mu\mathbf{d}_{ia}=\mu(\mathbf{d}_{i}-\mathbf{d}_{a}),$ (10c)
where $\mathbf{d}_{a}$ is the position of the target location (attraction point). The force $\mathbf{f}_{mi}$ of Eq. (10c) also points away from the target location $\mathbf{d}_{a}$ and pulls the cell towards the attraction point. The difference between push and pull forces is that the push force is stronger the further the distance to $\mathbf{d}_{m}$ is (and therefore the shorter the distance to $\mathbf{d}_{a}$ is), whereas the pull force is weaker the shorter the distance to $\mathbf{d}_{a}$ is. The push force provides a more efficient attraction to the target point at the cost of a greater dispersion of motile cells. From now on, we use the push force as our migration force of choice.
Fig 3: Simulated migration force towards an attraction point on the upper right corner. Migration of motile blue cells ($E_{\theta m}=15$ kPa) from passive red cell aggregate ($E_{\theta c}=1$ kPa), $j_{mm}=j_{cm}=j_{cc}=2$, $j_{ms}=j_{cs}=1$. Substrate strains are shown by black arrows. With Fig 1(a) as initial condition, result at 1500 MCTS of: (a) choice A, (b) choice B, (c) choice B with $j_{cc}=j_{cs}=1$, $j_{mm}=6$, $j_{cm}=4$, $j_{ms}=1/4$. See Video S2 suppl .
IV.0.1 Performance of the migration force
Incorporating the active push migration force to the CPM means solving Eqs. (10) and then implementing the Metropolis algorithm according to Eqs. (2)-(9). The strain gradient created by the migration force directs motile cells towards the point of attraction and away from the reference point. With respect to the YMT, we have two choices:
  • A.
    In extension (resp. retraction), adopt the YMT of the target (resp. source) pixel. This extends the extension-retraction symmetry of durotaxis under traction forces $\mathbf{f}$ to the combined force $\mathbf{f}+\mathbf{f}_{m}$.
  • B.
    Always adopt the YMT of the target pixel for $\mathbf{f}+\mathbf{f}_{m}$.
[…]
IV.0.2 Fractional step CPM